# Tracing the impacts of Mount Pinatubo eruption on global climate using spatially-varying changepoint detection

### v1

@@ Abstract
Significant events such as volcanic eruptions can have global and long lasting impacts on climate. These global impacts, however, are not uniform across space and time. Understanding how the Mt. Pinatubo eruption affects global and regional climate is of great interest for predicting impact on climate due to similar events. We propose a Bayesian framework to simultaneously detect and estimate spatially-varying temporal changepoints for regional climate impacts. Our approach takes into account the diffusing nature of the changes caused by the volcanic eruption and leverages spatial correlation. We illustrate our method on simulated datasets and compare it with an existing changepoint detection method. Finally, we apply our method on monthly stratospheric aerosol optical depth and surface temperature data from 1985 to 1995 to detect and estimate changepoints following the 1991 Mt. Pinatubo eruption.
Keywords Aerosol optical depth; Bayesian hierarchical model; Spatially-varying Changepoint, Temperature, Volcanic eruption
@@ 1 Introduction
The Mount Pinatubo eruption in June 1991 is the largest volcanic eruption in recent history that injected nearly 20 megatons of sulfur dioxide into the atmosphere. The aerosol cloud resulting from the eruption encircled the globe for a few weeks, and global changes to the atmosphere, including a decrease in surface temperature in the northern hemisphere of about 0.5 C, were observed up to two years after the event (Self et al., 1996). Studying the impact of volcanic eruptions are of high interest to the climate community since they serve as natural examples of Stratospheric Aerosol Injection (SAI), a controversial proposed solar climate intervention to help cool the earth and reduce the impact of global warming through a global dimming effect (Robock, 2014). Consequently, we aim for a more comprehensive understanding of Pinatubo to shed new light on this contentious technology against climate change.
@@
Scientists have studied how radiative forcing and climate were altered by the Mt. Pinatubo eruption through changes in aerosol optical depth. They found that the radiative forcing changes lasted for about 3 years (Stenchikov et al., 2009), leading to an observed heating in the stratosphere of 2-3K (Labitzke and McCormick, 1992) and cooling of global surface temperatures of about 0.4 K (Dutton and Christy, 1992; Thompson et al., 2009). Although there is a consensus that the eruption “caused" these global atmospheric changes, most published studies provided only qualitative evidence at a global scale. In this study, we attempt to quantitatively trace the impact of Pinatubo eruption on aerosol optical depth and temperatures at a finer scale.
@@
Since climate impacts from a volcanic eruption, or a general SAI event, are expected to cause sudden changes in regional climate. It is natural to use changepoints to trace the impact of eruption at regional scale. Changepoint detection methods aim to identify significant shifts in the underlying distribution of the data and have been employed to understand the impact of significant events on environment. For instance, Hallema et al. (2017) used changepoint analysis to assess the impacts of wildland fires on watershed annual water yield. Additionally, Robbins et al. (2011) developed a new changepoint test method to identify changes in the tropical cyclone record in the North Atlantic Basin along with climate change over the period 1851–2008.
@@
There have been many developments in changepoint detection in both univariate or multivariate settings, including two review articles (Reeves et al., 2007; Aminikhanghahi and Cook, 2017) and references therein. Additionally, methods such as pruned exact linear time (PELT) algorithm (Killick et al., 2012) and product partition model (PPM) (Barry and Hartigan, 1992) have gained considerable popularity in the changepoint literature. In situations where time series are indexed by spatial locations, the change behavior of time series that are geographically nearby is likely to be similar due to spatial correlation. For such data, the direct application of traditional changepoint detection methods to each time series might overlook this spatial feature and fail to exploit spatial correlation for improved changepoint detection and estimation.
@@
To address this limitation, Majumdar et al. (2005) employed a Bayesian approach to detect various types of changepoints by incorporating spatial correlation. Additionally, Xuan and Murphy (2007) extended the PPM to account for the dependency structure in multivariate time series using sparse Gaussian graphical models. While these methods can handle spatial dependence among time series, they assume a common changepoint across all spatial locations, imposing a strong limitation on their application. Recently, a spatially varying changepoint model was proposed by Berchuck et al. (2019) to monitor glaucoma progression in visual field data, where multivariate conditional autoregressive priors are used to model the spatial dependencies in the changepoints and the mean parameters. However, their model assumes the data is temporally independent. Another related work is Wang et al. (2023), who developed a spatially varying changepoint estimation method for functional time series observed over a spatial domain.
@@
The changes in regional climate caused by Pinatubo eruption are expected to exhibit similarity at nearby locations due to spatial dependency. However, these changes are anticipated to occur at different times due to the diffusion of the injected aerosols across the globe. In particular, we expect that the time of the change will increase as the spatial distance from the event location increases, again due to the diffusion of aerosols. These unique characteristics of our data render existing methods inadequate for our application. Consequently, we propose a novel space-time changepoint detection method that can identify when, where, and to what extent climate impacts have occurred following the 1991 eruption of Mt. Pinatubo while respecting the particular features of the SAI event. Our approach departs from conventional methods by modeling changepoints as a spatial process, where the expected value increases with the spatial distance from the event origin. By considering the spatial correlation of the data and the expected diffusion of the observed impact from the event location, our method is expected to be more effective in capturing the spatial patterns of climate impacts associated with the SAI events. Since our focus is soley on tracing the impact of the Mt. Pinatubo eruption, we examine at most one change after the eruption.
Our method extends the work of Majumdar et al. (2005) by allowing the changepoints to vary spatially. An additional advantage of our method is that both detection and estimation of changepoint are integrated in our approach. While most traditional methods focus on the mean shift, our method can detect changes in either the mean or variance and thus offers greater flexibility.
@@
The rest of this paper is organized as follows. Section 2 describes the data and presents exploratory data analysis results. Section 3 introduces the model formulation and sampling procedure. In Section 4, we conduct extensive simulation studies to evaluate the performance of our proposed method. Section 5 demonstrates our methodology on global aerosol optical depth (AOD) and surface temperature data observed around the time of the Mt. Pinatubo eruption. Finally, in section 6 we provide a brief summary and discuss future potential work.
@@ 2 Data
As mentioned in the 1, notable changes in AOD and surface temperatures caused by the Pinatubo eruption has been documented. While aerosols rapidly spread around the globe and experienced a near immediate and steep increase in magnitude following the eruption, changes in surface temperature were more subtle and took years to manifest. We consider the AOD that measures the extinction optical thickness at 550nm, named “TOTEXTTAU” in Global Modeling and Assimilation Office (2015b) (GMAO), and the surface temperature, named “tavg1_2d_slv_Nx” in Global Modeling and Assimilation Office (2015a) (GMAO), to evaluate the impact of the eruption. We obtained both data sets from the Modern-Era Retrospective analysis for Research and Applications, Version 2 (MERRA-2) reanalysis.
@@
The monthly AOD data ranges from January 1985 to December 1995, covering 132 time points in total. The data is over a $48\times 24$ grid and covers the entire globe. Following (Self et al., 1996), we exclude latitudes above $60^{\circ}$ and below $60^{\circ}$ in our analysis, resulting in $16\times 48=768$ spatial locations. Figure 1(a) shows the AOD data on these 768 spatial grids. In general, the aerosol data exhibits a clear jump after the eruption, followed by an approximately linear decay in time toward its previous level. However, not all locations precisely agree on the time of the jump, likely due to the aerosol diffusion process. To enhance our analysis, we first remove the seasonality by applying Seasonal-Trend decomposition using LOESS (STL) (Cleveland, 1990) to AOD data at each location. Then, we estimate the temporal trend as a linear function of time using the historical data up to June 1991, the month of the eruption. If the estimated trend parameter is statistically significant at a 0.05 level, we remove the estimated trend for both pre- and post-eruption data, detrending the time series before the changepoint to ensure a mean of zero. Finally, we normalize the time series at each location to achieve approximately constant variance across all locations.
@@
Fig. 1: Time Series of (a) stratospheric AOD data and (b) surface temperature data before preprocessing. Each blue curve represents a time series at a grid point. The red curve represents the global average.
@@
The monthly surface temperature data, shown in Figure 1(b), also spans from January 1985 to December 1995 and is on the same $48\times 24$ grid as the AOD data. In contrast to the AOD data, the impact of volcanic eruption on temperatures is indiscernible through simple visualization, suggesting a more subtle influence on temperatures, if any. When examining how temperature responds to a volcanic eruption, it is customary to study the temperature zonal mean indexed by latitude (Robock and Mao, 1995; Stenchikov et al., 2002; Gao et al., 2008). Hence, we average temperatures over longitude for each of the 24 latitude bands given in the data and conduct the analysis on the latitudinal means. This reduces the number of observations to $24$, with each latitude representing a latitude band covering $7.5^{\circ}$. We employed the same procedure as the AOD data to detrend and normalize temperature data. The processed zonal mean temperatures can be seen in Figure 8.
@@
To evaluate our conjecture that changes in these climate series vary spatially, we conducted exploratory data analysis on the global AOD data. At each grid location, we applied the Bayesian information criteria-based chagepoint detection procedure, which is a well-established method for climate series (Reeves et al., 2007). Figure 2 illustrates the heatmap of detected changepoints and estimated mean shifts at each grid location. Overall, the changepoints appear to be spatially varying and clustered, as does the magnitude of change. Although a few locations show changepoints prior to 1991/06 that are apparently not due to Mt. Pinatubo eruption, in general, the changepoints appear to spread latitudinally from the source of the eruption. This spatial pattern is consistent with what we expect from an SAI event, given the influence of strong zonal winds in the lower stratosphere that circulate the globe (Robock and Matson, 1983). These results not only verify our conjecture but also provide guidance for the development of our methodology.
@@
Fig. 2: (a) Heatmap of detected changepoints for AOD. Darker color indicates earlier change. (b) Heatmap of estimated change amount. Darker color indicates stronger magnitude of mean shift. The red diamond marks the location of Mt. Pinatubo.
@@ 3 Method
We propose a Bayesian hierarchical model to detect and estimate spatially varying changepoints for a spatiotemporal data. We model the likelihood of data in the first level, provide the prior for latent processes in the second level, and close the hierarchy by specifying priors and hyperpriors for all unknown parameters in the model. We then discuss the challenges of MCMC for the proposed model.
@@ 3.1 Changepoint detection and estimation
Let $Y(\mathbf{s},t)$ denote the observed climate variable at location $\mathbf{s}\in\mathbb{R}^{2}$ and time $t\in\mathbb{N}$, $N$ be the total number of spatial locations, and $M$ the total time points. We propose the following model for the spatiotemporal process $Y(\mathbf{s},t)$ with possible mean and variance change in time:
@@
$Y(\mathbf{s},t)$ =$=\begin{cases}\mu_{1}(\mathbf{s},t)+U(\mathbf{s},t)+\epsilon_{1}(\mathbf{s},t),&t\leq\tau(\mathbf{s})\\ \mu_{2}(\mathbf{s},t)+U(\mathbf{s},t)+\epsilon_{2}(\mathbf{s},t),&t>\tau(\mathbf{s}),\end{cases}$ (1)
where $\tau(\mathbf{s})$ is the changepoint at location ${\bf s}$, $\mu_{1}(\mathbf{s},t)$ and $\mu_{2}(\mathbf{s},t)$ are the mean functions before and after changepoint, $U(\mathbf{s},t)$ is a zero mean space-time correlated error process, and $\epsilon_{1}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma_{1}^{2})$ and $\epsilon_{2}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma_{2}^{2})$ are the before and after changepoint measurement errors. When $\tau(\mathbf{s})=M$, it indicates the absence of changepoint at ${\bf s}$ as the entire process is governed by a single model. To keep our model parsimonious and avoid potential identifiablity issue, we assume a common spatiotemporal process $U(s,t)$ for both before and after changepoint. This ensures that all changes will be attributed to mean and/or variance of measurement errors. This assumption also allows us to take advantage of the space-time separablility of the covariance structure to facilitate the computation for spatiotemporal data. While Majumdar et al. (2005) allows for a change in the covariance structure by modeling the pre-changepoint error process as $U({\bf s},t)+V({\bf s},t)$ and the post-changepoint error process as $U({\bf s},t)+W({\bf s},t)$, their model assumes $\tau({\bf s})$ as well as $\mu({\bf s},t)$ to be a constant as a sacrifice to reduce the identifiability risk. Since the primary interest for our data is to identify the spatially varying changepoints while respecting the spatio-temporally varying mean, we opt for a model that provides this flexibility.
As mentioned in Section 2, the data after preprocessing can be assumed to have constant mean prior to the changepoint and a linear trend in time after changepoint. We also observed from Figure 2 that the change amount not only varies spatially but also tends to be clustered in space. To account for these features, we model the pre and post-changepoint mean processes as
@@
In (2) and (3), $\alpha_{0}$ is the constant parameter representing the average global mean before the changepoint, $\gamma_{0}({\bf s})$ is the mean shift at the time of the changepoint at location ${\bf s}$, and $\gamma_{1}({\bf s})$ represents the slope after changepoint at locations ${\bf s}$. We further decompose $\gamma_{0}({\bf s})$ and $\gamma_{1}({\bf s})$ into a fixed and random components in (4). The parameters $\gamma_{0F}$ and $\gamma_{1F}$ can be interpreted as the average mean shift and post-changepoint temporal trend across all locations that have a changepoint. The parameters $\gamma_{0R}({\bf s})$ and ${\gamma_{1R}({\bf s})}$ serve as spatial random effects to allow the mean shift and trend to vary spatially. To respect the spatial correlation of change amount and the post-changepoint trend, we model $\gamma_{0R}({\bf s})$ and ${\gamma_{1R}({\bf s})}$ as a spatially correlated process.
@@
Let ${\bf s}_{0}$ denote the location at which the change is first observed, i.e., ${\bf s}_{0}=\arg\min_{\bf s}\tau({\bf s}).$ Assuming that the change diffuses from the event origin, ${\bf s}_{0}$ can also be interpreted as the event location, i.e., Mt. Pinatubo in our context. We model $\tau(\mathbf{s})$ as the sum of two terms: (1) the changepoint at the event location ${\bf s}_{0}$, denoted by $\tau_{0}$, and (2) the temporal lag between changepoints at $\mathbf{s_{0}}$ and ${\bf s}$, denoted by $\Delta(\mathbf{s})$. Specifically, we have
$\tau(\mathbf{s})=\min\{M,\tau_{0}+\Delta(\mathbf{s})\}.$ (5)
@@
We require $\Delta(\mathbf{s})\geq 0$ such that $\tau_{0}$ is the earliest changepoint, which ensures that all other changepoints detected are at or after $\tau_{0}$ and thus not due to other unrelated events. We also fix $\Delta(\mathbf{s}_{0})=0$ in our estimation so that $\tau(\mathbf{s}_{0})=\tau_{0}$ is observed as the earliest changepoint, i.e., $\tau(\mathbf{s}_{0})=\min_{{\bf s}}\tau({\bf s})$. Note that $\tau_{0}$ takes discrete values in $\{1,\ldots,M\}$ and $\tau_{0}=M$ implies that there is no changepoint at any of the locations. There are two main advantages of separating $\tau_{0}$ from $\tau(\mathbf{s})$. First, it ensures that the changepoint at any location $\mathbf{s}\neq\mathbf{s}_{0}$ occurs strictly after the changepoint at $\mathbf{s}_{0}$, as any changepoint prior to $\tau_{0}$ would be due to another event that is not of our interest. Second, we can model $\Delta(\mathbf{s})$ as a continuous process with a spatial correlation structure to reflect the spatial pattern observed in Figure 2(a) while keeping $\tau_{0}$ as a discrete variable. This allows for a convenient way to test for the existence of the changepoint and also incorporate the prior knowledge about the possible range of date at which the change was first observed from the eruption.
@@
We assume $\boldsymbol{\Delta}=(\Delta({\bf s}_{1}),\ldots,\Delta({\bf s}_{N}))^{T}$ follow a log normal process with mean $\mathbf{X}\boldsymbol{\beta}$ and covariance matrix ${\bf\Sigma}_{\Delta}$, where $\mathbf{X}$ is the $N\times 2$ matrix of spatial lags from each location $\mathbf{s}$ to the event location ${\bf s}_{0}$, with the $i$th row of $\boldsymbol{X}$ given by $(\hbox{lon}(\mathbf{s}_{i})-\hbox{lon}(\mathbf{s}_{0}),\hbox{lat}(\mathbf{s}_{i})-\hbox{lat}(\mathbf{s}_{0}))$. $\boldsymbol{\beta}$ is the 2-dimensional vector with each component representing the rate of diffusion in longitude and latitude, respectively.
In other words, we model $\Delta({\bf s})$ as a spatial process that increases with the distance from $s_{0}$ to reflect the nature of the physical process we observe, i.e. the time it takes for the impact of the event to reach a specific location will tend to increase as its distance from the event location increases.

### v2

@@ Abstract
Significant events such as volcanic eruptions can have global and long-lasting impacts on climate. These global impacts, however, are not uniform across space and time. Understanding how the Mt. Pinatubo eruption affects global and regional climate is of great interest for predicting the impact on climate due to similar events as well as understanding the possible effect of the Stratospheric Aerosol Injections proposed to combat climate change. While many studies illustrated the impact of the Pinatubo eruption on a global scale, studies at a fine regional scale are scarce. We propose a Bayesian spatially-varying changepoint detection and estimation method to trace the impact of Mt Pinatubo eruption on regional climate. Our approach takes into account the diffusing nature and spatial correlation of the climate changes attributed to the volcanic eruption. We illustrate our method and demonstrate its advantages over an existing changepoint detection method through simulations. Finally, we apply our method to monthly stratospheric aerosol optical depth and surface temperature data from 1985 to 1995 to detect and estimate changepoints following the 1991 Mt. Pinatubo eruption. Our results quantitatively characterize the spatial pattern of the eruption’s impact on regional climate, complementing the previous studies on the global impact of the Pinatubo eruption.
Keywords Aerosol optical depth, Bayesian hierarchical model, Spatially-varying Changepoint, Temperature, Volcanic eruption
@@ 1 Introduction
The Mount Pinatubo eruption in June 1991 is the largest volcanic eruption in recent history that injected nearly 20 megatons of sulfur dioxide into the atmosphere. The aerosol cloud resulting from the eruption encircled the globe for a few weeks, and global changes to the atmosphere, including a decrease in surface temperature in the northern hemisphere of about 0.5 C, were observed up to two years after the event (Self et al., 1996). Studying the impact of volcanic eruptions is of high interest to the climate community for two main reasons. On the one hand, it helps improve predictions of the anticipated impacts from new volcanic activities or other localized events such as wildfires. On the other hand, volcanic eruptions serve as natural examples of Stratospheric Aerosol Injection (SAI), a controversial proposed solar climate intervention to help cool the earth and reduce the impact of global warming through a global dimming effect (Robock, 2014). Thus, a more comprehensive understanding of the impact of the Pinatubo eruption can shed new light on this radical technology against climate change.
@@
Scientists have extensively studied how atmospheric properties were altered by the Mt. Pinatubo eruption. They found that the radiative forcing changes lasted for about three years (Stenchikov et al., 2009), leading to observed heating in the stratosphere of 2-3K (Labitzke and McCormick, 1992) and cooling of global surface temperatures of about 0.4 K (Dutton and Christy, 1992; Thompson et al., 2009). Although there is a consensus that the eruption “caused" these global atmospheric changes, most published studies provided only qualitative evidence on a global scale. In this study, we attempt to quantitatively trace the impact of the Pinatubo eruption on aerosol optical depth and temperatures at a finer scale.
@@
Since climate impacts from a volcanic eruption or a general SAI event are expected to cause sudden changes in regional climate, it is natural to use changepoints to trace the impact of an eruption at a regional scale. Changepoint detection methods aim to identify significant shifts in the underlying distribution of the data and have been employed to understand the impact of significant events on the environment. For instance, Hallema et al. (2017) used changepoint analysis to assess the impacts of wildland fires on watershed annual water yield, Robbins et al. (2011) developed a new changepoint detection method to identify changes in the tropical cyclone record in the North Atlantic Basin along with climate change over the period 1851–2008, and Tucker and Yarger (2023) found evidence of increased stratospheric temperature after the Mt. Pinatubo eruption using an epidemic changepoint model.
@@
There have been many developments in changepoint detection in both univariate and multivariate settings, including two review articles (Reeves et al., 2007; Aminikhanghahi and Cook, 2017) and references therein. In particular, methods such as pruned exact linear time (PELT) algorithm (Killick et al., 2012) and product partition model (PPM) (Barry and Hartigan, 1992) have gained considerable popularity in the changepoint literature. In situations where time series are indexed by spatial locations, the change behavior of time series that are geographically nearby is likely to be similar due to spatial correlation. For such data, the direct application of traditional changepoint detection methods to each time series might overlook this spatial feature and is inadequate to exploit spatial correlation to improve changepoint detection and estimation.
@@
To address this limitation, Majumdar et al. (2005) employed a Bayesian approach to detect various types of changepoints by incorporating spatial correlation. Additionally, Xuan and Murphy (2007) extended the PPM to account for the dependency structure in multivariate time series using sparse Gaussian graphical models. While these methods can handle spatial dependence among time series, they assume a common changepoint across all spatial locations, imposing a strong limitation on their application. Recently, a spatially-varying changepoint model was proposed by Berchuck et al. (2019) to monitor glaucoma progression in visual field data, where multivariate conditional autoregressive priors are used to model the spatial dependencies in the changepoints and the mean parameters. Their model assumed the data is temporally independent. Another related work is Wang et al. (2023), who developed a spatially-varying changepoint estimation method for functional time series observed over a spatial domain.
@@
The changes in regional climate caused by the Pinatubo eruption or a general SAI event are expected to exhibit similarity at nearby locations due to spatial dependency. However, these changes are anticipated to occur at different times due to the diffusion of the injected aerosols across the globe and the atmospheric circulation. For particular variables such as the global aerosol optical depth (AOD) that measures the extinction of the solar beam by dust and haze, we expect that the time of the change will increase as the spatial distance from the event location increases, again due to the diffusion of aerosols. These unique characteristics of our problem render existing methods inadequate for our application. We, therefore, propose a novel space-time changepoint detection method that can identify when, where, and to what extent climate impacts have occurred following the 1991 eruption of Mt. Pinatubo while respecting the particular features of the SAI event.
Our approach departs from conventional methods by modeling changepoints as a spatial process and further allowing time-after-event of the changepoints to increase with the spatial distance from the event origin. By taking into account the spatial correlation of the data and the anticipated diffusion of the observed impact from the event location, our method is demonstrated to be more effective in capturing the spatial patterns of climate impacts associated with the SAI events. Since our focus is solely on tracing the impact of the Mt. Pinatubo eruption, we consider at most one change after the eruption. Compared to Majumdar et al. (2005), our method extends their model by allowing for spatially-varying changepoints. Furthermore, our approach performs both detection and estimation of changepoint at each location rather than only the estimation. Lastly, while most traditional methods focus on the mean shift, our method can detect changes in either the mean or variance, thus offering greater flexibility.
@@
The rest of this paper is organized as follows. Section 2 describes the data and presents exploratory data analysis results. Section 3 introduces the model formulation and sampling procedure. In Section 4, we conduct extensive simulation studies to evaluate the performance of our proposed method. Section 5 demonstrates our methodology on AOD and surface temperature data observed around the time of the Mt. Pinatubo eruption. Finally, in Section 6 we provide a brief summary and discuss potential future work.
@@ 2 Data
As mentioned in Section 1, notable changes in aerosol optical depth (AOD) and surface temperatures caused by the Pinatubo eruption have been documented. While aerosols rapidly spread around the globe and experienced a near immediate and steep increase in magnitude following the eruption, changes in surface temperature were more subtle and took years to manifest. We consider the stratospheric AOD that measures the extinction optical thickness at 550nm111referred to as“TOTEXTTAU” in Global Modeling and Assimilation Office (2015a) (GMAO), and the surface temperature222referred to as “tavg1_2d_slv_Nx” in Global Modeling and Assimilation Office (2015b) (GMAO), to evaluate the impact of the eruption. We obtained both data sets from the Modern-Era Retrospective analysis for Research and Applications, Version 2 (MERRA-2) [^3] https://gmao.gsfc.nasa.gov/reanalysis/MERRA-2/.
@@
The monthly AOD data ranges from January 1985 to December 1995, covering 132 time points in total. The data is over a $48\times 24$ grid and covers the entire globe. Following Self et al. (1996), we exclude latitudes above $60^{\circ}$N and below $60^{\circ}$S in our analysis, resulting in $48\times 16=768$ spatial locations. Figure 1(a) shows the AOD data on these 768 spatial grids. In general, the aerosol data exhibits a clear jump after the eruption, followed by an approximately linear decay in time toward its previous level. However, not all locations precisely agree on the jump time, likely due to the aerosol diffusion process. To enhance our analysis, we first remove the seasonality by applying Seasonal-Trend decomposition using LOESS (STL) (Cleveland, 1990) to AOD data at each location. Then, we estimate the temporal trend as a linear function of time using the historical data up to June 1991, the month of the eruption. If the estimated trend parameter is statistically significant at a 0.05 level, we remove the estimated trend for both pre- and post-eruption data. Assuming the trend between the month of eruption and the actual changepoint remains the same as in the pre-eruption period, this procedure roughly centers the time series before the changepoint so a constant mean can be assumed. Finally, we normalize the time series at each location by calculating the sample variance using data before June 1991 and use this number to achieve approximately constant variance across all locations.
@@
Fig. 1: Time Series of (a) stratospheric AOD data and (b) surface temperature data before preprocessing. Each grey curve represents a time series at a grid point. The black curve represents the global average.
@@
The monthly surface temperature data, shown in Figure 1(b), also spans from January 1985 to December 1995 and is on the same $48\times 24$ grid as the AOD data. In contrast to the AOD data, the impact of volcanic eruption on temperatures is indiscernible through simple visualization, suggesting a more subtle influence on temperatures, if any. When examining how temperature responds to a volcanic eruption, it is customary to study the temperature zonal mean indexed by latitude (Robock and Mao, 1995; Stenchikov et al., 2002; Gao et al., 2008). Hence, we average temperatures over longitude for each of the 24 latitude bands given in the data and conduct the analysis on the latitudinal means. This reduces the number of observations to $24$, with each latitude representing a latitude band covering $7.5^{\circ}$. We employed the same procedure as for the AOD data to detrend and normalize temperature data. The processed zonal mean temperatures can be seen in Figure 8.
@@
To evaluate our conjecture that changes in these climate variables vary spatially, we conducted exploratory data analysis on the global AOD data. At each grid location, we applied the Bayesian information criteria-based changepoint detection procedure, which is a well-established method for climate series (Reeves et al., 2007). Figure 2 illustrates the heatmap of detected changepoints and estimated mean shifts at each grid location. Overall, the changepoints appear to be spatially-varying and clustered, as does the magnitude of change. Although a few locations show changepoints prior to 1991/06 that are apparently not due to the Mt. Pinatubo eruption, in general, the changepoints appear to spread latitudinally from the source of the eruption. This spatial pattern is consistent with what we expect from an eruption event, given the influence of strong zonal winds in the lower stratosphere that circulate the globe (Robock and Matson, 1983). These results not only verify our conjecture but also provide guidance for the development of our methodology.
@@
Fig. 2: (a) Heatmap of detected changepoints for AOD. Darker color indicates earlier change. (b) Heatmap of estimated change amount. Darker color indicates stronger magnitude of mean shift. The white diamond marks the location of Mt. Pinatubo.
@@ 3 Method
We propose a Bayesian hierarchical model to detect and estimate spatially-varying changepoints for spatiotemporal data. We model the likelihood of the data in the first level, provide the prior for the latent processes in the second level, and close the hierarchy by specifying priors and hyperpriors for all unknown parameters in the third level. We then discuss the challenges and the schemes of the Markov Chain Monte Carlo (MCMC) sampling algorithm used to estimate the proposed model.
@@ 3.1 Changepoint detection and estimation
Let $Y(\mathbf{s},t)$ denote the observed climate variable at location $\mathbf{s}\in\mathbb{S}^{2}$ and time $t\in\mathbb{N}$, $N$ be the total number of spatial locations, and $M$ the total time points. We propose the following model for the spatiotemporal process $Y(\mathbf{s},t)$ with possible mean and variance changes in time:
@@
$Y(\mathbf{s},t)$ $=\begin{cases}\mu_{1}(\mathbf{s},t)+U(\mathbf{s},t)+\epsilon_{1}(\mathbf{s},t),&t\leq\tau(\mathbf{s})\\ \mu_{2}(\mathbf{s},t)+U(\mathbf{s},t)+\epsilon_{2}(\mathbf{s},t),&t>\tau(\mathbf{s}),\end{cases}$ (1)
where $\tau(\mathbf{s})$ is the changepoint at location $\bf s$}, $\mu_{1}(\mathbf{s},t)$ and $\mu_{2}(\mathbf{s},t)$ are the mean functions before and after changepoint, $U(\mathbf{s},t)$ is a zero mean space-time correlated error process, and $\epsilon_{1}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma_{1}^{2})$ and $\epsilon_{2}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}N(0,\sigma_{2}^{2})$ are the before and after changepoint measurement errors. When $\tau(\mathbf{s})=M$, it indicates the absence of changepoint at location $\bf s$} as the entire process is governed by a single model. To keep our model parsimonious and avoid potential identifiability issues, we assume a common spatiotemporal process $U(s,t)$ both before and after changepoint. This ensures that all changes will be attributed to mean and/or variance of measurement errors. This assumption also allows us to take advantage of the possible space-time separable covariance structure to facilitate the computation for spatiotemporal data. While Majumdar et al. (2005) allows for a change in the covariance structure by modeling the pre-changepoint error process as $U({\bf s},t)+V({\bf s},t)$ and the post-changepoint error process as $U({\bf s},t)+W({\bf s},t)$, their model assumes $\tau({\bf s})$ as well as $\mu({\bf s},t)$ to be constant as a sacrifice to ensure identifiability. Since the primary interest for our data is to identify the spatially-varying changepoints while respecting the spatiotemporally-varying mean, we opt for a model that offers this flexibility.
As mentioned in Section 2, the data after preprocessing can be assumed to have a constant mean prior to the changepoint. Based on Figure 1(a), we assume a linear trend in time after changepoint. We also observe from Figure 2 that the change amount not only varies spatially but also tends to be clustered in space. To account for these features, we model the pre and post-changepoint mean processes, $\mu_{1}$ and $\mu_{2}$ respectively, as
@@
In (2) and (3), $\alpha_{0}$ is the constant parameter representing the average global mean before the changepoint, $\gamma_{0}({\bf s})$ is the mean shift at the time of the changepoint at location $\bf s$}, and $\gamma_{1}({\bf s})$ represents the slope after changepoint at locations $\bf s$}. We further decompose $\gamma_{0}({\bf s})$ and $\gamma_{1}({\bf s})$ into a fixed and random components in (4). The fixed components $\gamma_{0F}$ and $\gamma_{1F}$ can be interpreted as the average mean shift and the average post-changepoint temporal trend across all locations that have a changepoint. The parameters $\gamma_{0R}({\bf s})$ and $\gamma_{1R}({\bf s})$} serve as spatial random effects to allow the mean shift and trend to vary spatially. To respect the spatial correlation of the magnitude of change and of the post-changepoint trend, we model $\gamma_{0R}({\bf s})$ and $\gamma_{1R}({\bf s})$} as a spatially correlated process.
@@
Let ${\bf s}_{0}$ denote the location at which the change is first observed among all locations where we attempt to detect the changepoint, i.e., ${\bf s}_{0}=\arg\min_{\bf s}\tau({\bf s})$. If the influence of the volcanic eruption on the variable of interest diffuses from the event origin, ${\bf s}_{0}$ is usually either the event origin or some location nearby depending on what locations we consider for changepoint detection. We model $\tau(\mathbf{s})$ as the sum of two terms: the changepoint at the location ${\bf s}_{0}$, denoted by $\tau_{0}$, and the temporal lag between changepoints at $\mathbf{s_{0}}$ and $\bf s$}, denoted by $\Delta(\mathbf{s})$. Specifically, we have
$\tau(\mathbf{s})=\min\{M,\lfloor\tau_{0}+\Delta(\mathbf{s})\rfloor\}.$ (5)
@@
In theory, the “true” changepoint process can be defined as $\tau_{0}+\Delta({\bf s})$, where the temporal lags $\Delta({\bf s})$ are continuous and take values in $(0,\infty)$. However, the observed changepoint is discrete and we are mainly interested in changepoints between 1 to M, so we take the floor sign and cap the value of changepoints at $M$. We require $\Delta(\mathbf{s})\geq 0$ such that $\tau_{0}$ is the earliest changepoint. This ensures that all other changepoints detected are either at or after $\tau_{0}$ and thus not absolutely due to other unrelated events. We fix $\Delta(\mathbf{s}_{0})=0$ in our estimation because $\tau(\mathbf{s}_{0})=\tau_{0}=\min_{{\bf s}}\tau({\bf s})$, as defined. Note that $\tau_{0}$ takes discrete values in $\{1,\ldots,M\}$ and $\tau_{0}=M$ implies that there is no changepoint at any of the locations. There are two main advantages of separating $\tau_{0}$ from $\tau(\mathbf{s})$. First, it ensures that the changepoint at any location $\mathbf{s}\neq\mathbf{s}_{0}$ occurs strictly after the changepoint at $\mathbf{s}_{0}$, as any changepoint prior to $\tau_{0}$ would be due to another event that is not of our interest. Second, we can model $\Delta(\mathbf{s})$ as a continuous process with a spatial correlation structure to reflect the spatial pattern observed in Figure 2(a) while keeping $\tau_{0}$ as a discrete variable. The discrete parameter $\tau_{0}$ can help us easily test if it is the case that no changepoint ever occurs at any of the locations and incorporate prior knowledge about the possible range of date at which the change caused by the eruption was first observed.
@@
We further assume $\boldsymbol{\Delta}=(\Delta({\bf s}_{1}),\ldots,\Delta({\bf s}_{N}))^{T}$ follow a log normal process with mean $\mathbf{X}\boldsymbol{\beta}$ and covariance matrix ${\bf\Sigma}_{\Delta}$, where $\mathbf{X}$ is the $N\times 2$ matrix of spatial lags from each location $\mathbf{s}$ to the location ${\bf s}_{0}$, with the $i$th row of $\boldsymbol{X}$ given by $(\hbox{lon}(\mathbf{s}_{i})-\hbox{lon}(\mathbf{s}_{0}),\hbox{lat}(\mathbf{s}_{i})-\hbox{lat}(\mathbf{s}_{0}))$, and $\boldsymbol{\beta}$ is the 2-dimensional vector with each component representing the rate of diffusion in longitude and latitude, respectively. Allowing $\Delta(\mathbf{s})$ to increase with the distance from ${\bf s}_{0}$ in an anisotropic manner reflects the nature of the physical process shown in Figure 2 that the eruption diffuses at a different rate along the longitude and latitude, but in any direction, it takes a longer time for the eruption impact to reach a specific location that is further away from the event origin.